Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $0.75 coming 3 years from today. The dividend should grow rapidly-at a rate of 22% per year-during Years 4 and 5; but after Year 5, growth should be a constant 9% per year. If the required return on Computech is 17%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

D3=0.75

D4=(0.75*1.22)=0.915

D5=(0.915*1.22)=1.1163

Value after year 5=(D5*Growth rate)/(Required rate-Growth rate)

=(1.1163*1.09)/(0.17-0.09)

=15.2095875

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=0.75/1.17^3+0.915/1.17^4+1.1163/1.17^5+15.2095875/1.17^5

=$8.40(Approx).